### v1

# TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling

Abstract
Large Language Models (LLMs) rely on attention mechanisms whose time complexity grows quadratically with input sequence length, creating significant computational bottlenecks during the prefilling stage. Existing static sparse attention methods typically degrade accuracy, while dynamic sparsity methods introduce additional computational overhead due to runtime sparse index estimation. To address these limitations, we propose TriangleMix, a novel training-free static attention pattern. TriangleMix employs dense attention in shallow layers and switches to a triangle-shaped sparse pattern in deeper layers. Extensive experiments demonstrate that TriangleMix reduces attention overhead by 3.7×3.7$\times$3.7 × to 15.3×15.3$\times$15.3 × in deep layers, and decreases overall Time-to-First-Token (TTFT) by 12% to 32% for sequence lengths ranging from 32K to 128K, without sacrificing model accuracy. Moreover, TriangleMix can be seamlessly integrated with dynamic sparsity methods to achieve further speedup, e.g. accelerating MInference by 19% at 128K, highlighting its potential to enhance LLM inference efficiency. Implementation at https://aka.ms/TriangleMix
TriangleMix: A Lossless and Efficient Attention Pattern for Long Context Prefilling
Zhiyuan He1, Yike Zhang2thanks: Work during internship at Microsoft. , Chengruidong Zhang1, Huiqiang Jiang1, Yuqing Yang1, Lili Qiu1 1Microsoft Research, 2Tsinghua University zhiyuhe@microsoft.com
1 Introduction
(a) TriangleMix
(b) TriangleMix combined with dynamic attention
Figure 1: TriangleMix on Llama-3.1-8B-Instruct.
Large Language Models (LLMs) are capable of processing input sequences of varying lengths Grattafiori et al. (2024); Achiam et al. (2023); Yang et al. (2024), a crucial ability that supports diverse downstream tasks, such as question answering Bai et al. (2023), long-form document understanding Zhao et al. (2024), and code generation Jiang et al. (2024b). However, due to the quadratic complexity of attention mechanisms, the time-to-first-token (TTFT) significantly increases as the input context length grows. As prior research has demonstrated, attention computation has become a critical bottleneck in the prefilling stage of LLMs Jiang et al. (2024a); Lai et al. (2025).
To address this bottleneck and accelerate the prefilling stage, researchers have proposed both static and dynamic sparse attention methods. Static sparse attention methods, such as StreamingLLM Xiao et al. (2023), reduce computational complexity from $O(N^{2})$ to $O(N)$ but suffer notable performance degradation on tasks with long contexts Li et al. (2024a). On the other hand, dynamic sparsity approaches, such as MInference Jiang et al. (2024a) and FlexPrefill Lai et al. (2025), aim to maintain near-full performance by dynamically computing sparse attention indices during inference.
Dynamic sparsity methods are effective for handling extremely long contexts (e.g., beyond 128K tokens). However, they incur additional overhead due to the runtime estimation of sparse block indices. While this overhead is negligible at very long sequence lengths, it becomes non-trivial for moderately long contexts ranging from 32K to 128K tokens. We measured the average attention kernel time per layer on Llama-3.1-8B-Instruct, and the results (Table 1) show that MInference can be 2× slower than dense attention at 32K, and offers only limited speedup at 64K. Similarly, FlexPrefill delivers a modest 1.4× speedup at 32K.
Figure 2: The average gradient $\mathrm{Grad}(\bm{M},l)$ of the Middle Q-K sections, measured on Llama-3.1-8B-Instruct, shows a significant decline in deeper layers. This suggests that the Middle Q-K components in deeper layers contribute minimally and might potentially be skipped to improve efficiency.
In this work, we aim to develop a static attention pattern that effectively accelerates the prefilling stage across various context lengths. To achieve this goal, we propose a novel gradient-based method to identify the crucial sections within attention maps. Our analysis demonstrates that the middle Q-K section (Figure 2 and 3) of the attention map can be safely omitted in deeper layers without incurring performance loss. Motivated by this insight, we introduce TriangleMix, a lossless and efficient static attention pattern for accelerating long-context LLMs. As depicted in Figure 1, TriangleMix applies standard dense attention in shallow layers, and switches to a triangle-shaped sparse attention pattern in deeper layers. TriangleMix offers four major advantages:
  • It is training-free, enabling direct integration with state-of-the-art pretrained LLMs.
  • It achieves nearly lossless accuracy, preserving the full capabilities of the underlying models.
  • Its static triangle-shaped sparsity eliminates the need for dynamic block index estimation and reduces the computational complexity from $O(N^{2})$ to $O(N)$.
  • It is complementary to dynamic attention approaches. Replacing dynamic attention in deep layers with the static Triangle attention further enhances acceleration, compared with dynamic sparse attention alone.
We conduct extensive experiments on three long-context LLMs, including Llama-3.1-8B-Instruct Grattafiori et al. (2024), Llama-3-8B-Instruct-262K GradientAI (2024), and Qwen2.5-7B-Instruct Yang et al. (2024), using all tasks from the RULER and LongBench benchmarks Hsieh et al. (2024); Bai et al. (2023). The results demonstrate that TriangleMix maintains the same accuracy as full attention. For Llama-3.1-8B-Instruct, TriangleMix decreases overall TTFT by 12% to 32% across context lengths ranging from 32K to 128K. Moreover, it can be seamlessly combined with dynamic sparsity methods without degrading performance. For example, integrating TriangleMix with MInference decreases TTFT by 19% at 128K context length for Llama-3.1-8B-Instruct.
2 Methodology
2.1 Sparse Attention
Figure 3: Three attention sections.
Figure 4: First row: Average attention score $\mathrm{Att}(\bm{M},l)$ for the Middle and Last Q-K sections; Second row: average gradient $\mathrm{Grad}(\bm{M},l)$ for the Middle and Last Q-K sections.
Method 32K 64K 128K
Dense 44 179 750
MInference 94 152 229
FlexPrefill ($\gamma=0.95$) 32 71 220
Table 1: Averaged attention kernel time in milliseconds for one layer on Llama-3.1-8B-Instruct.
The prefilling stage of Transformer attention can be formulated as:
$\bm{A}=\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M}))$
[…]
To accelerate computing, sparse attention aims to find a sparse mask matrix $\bm{M}^{\prime}$ to compute the attention output:
$\bm{A^{\prime}}=\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M^{\prime}}))$
where $|\bm{A}-\bm{A}^{\prime}|$ is expected to be sufficiently small to preserve the original model performance.
Existing sparse attention methods fall into two categories: static sparsity and dynamic sparsity. In static sparsity, the sparse mask $\bm{M}^{\prime}$ is fixed and does not depend on the input. For example, StreamingLLM shows that most attention scores accumulate on the first few sink tokens and nearby tokens in a sliding window Xiao et al. (2023). This leads to the following streaming pattern:
$\bm{M}^{\mathrm{streaming}}_{i,j}=\begin{cases}1,&i\geq j,j\leq si\\ 1,&i\geq j,i-j\leq sl\\ 0,&\mathrm{otherwise}\end{cases}$
where $si$ is the number of sink tokens and $sl$ is the sliding window size. This pattern, also called "A-shape" in previous works Jiang et al. (2024a), is very efficient with only $O(N)$ complexity. However, it causes significant performance drops for long-context LLMs Li et al. (2024a). In contrast, dynamic sparsity methods, such as MInference Jiang et al. (2024a) and FlexPrefill Lai et al. (2025), estimate $\bm{M}^{\prime}$ at inference time based on the input through an observation window or the pooling vector of $\bm{Q}$ and $\bm{K}$.
Dynamic sparsity generally provides better accuracy and acceleration for extremely long contexts (e.g., contexts exceeding 128K tokens). However, for moderately long contexts, the online estimation introduces additional time overhead. Specifically, as demonstrated in Table 1, MInference is slower than dense attention or offers small acceleration for contexts below 64K, while FlexPrefill achieves only modest improvements (approximately 1.4x speedup for 32K context).
2.2 Attention Sections
In this paper, we aim to design a static sparse attention mask that is both efficient and effective. Although a large fraction of attention scores falls into the streaming pattern, a significant performance gap still remains between streaming sparse attention and full dense attention. This leads us to the following question:
Beyond the streaming pattern, what other parts of the attention are crucial for maintaining the end-to-end performance of long-context LLMs?
To answer this, we divide the full attention matrix into three distinct sections as illustrated in Figure 3:
  • Streaming section: includes attention sink and the sliding window;
  • Last Q-K section: covers interactions between the last part of $\bm{Q}$ and $\bm{K}$, excluding the Streaming section;
  • Middle Q-K section: consists of the remaining interactions between the middle parts of $\bm{Q}$ and $\bm{K}$.
We define the Last Q-K mask as:
$\bm{M}^{\mathrm{last}}_{i,j}=\begin{cases}1,&i\geq j,N-i<last,\\ &j>si,i-j>sl\\ 0,&\mathrm{otherwise}\end{cases}$
[…]
Similarly, the Middle Q-K mask is defined as:
$\bm{M}^{\mathrm{middle}}_{i,j}=\begin{cases}1,&i\geq j,N-i\geq last,\\ &j>si,i-j>sl\\ 0,&\mathrm{otherwise}\end{cases}$
2.3 Probing Attention Importance
Firstly, we inspect the average attention scores within the two sections defined by masks $\bm{M}^{\mathrm{middle}}$ and $\bm{M}^{\mathrm{last}}$. We observe that these average attention scores are quite similar, although both are notably lower than the average score observed in the streaming section $\bm{M}^{\mathrm{streaming}}$. To achieve a more precise measurement of the relative importance of these sections, we introduce a novel gradient-based probing technique.
Specifically, given an input $\bm{X}_{\mathrm{input}}$ containing $N$ tokens fed into a language model, we define a probing variable $\bm{\theta}\in\mathbb{R}^{L\times N\times N}$, where $\bm{\theta}_{\ell,i,j}=1$ for all layers $\ell$ and token indices $i,j$. Here, $L$ represents the total number of layers in the model. For simplicity of notation, we omit the attention head dimension here, but our proposed method generalizes naturally to both Multi-head Attention Vaswani et al. (2017) and Grouped Query Attention Ainslie et al. (2023).
In layer $l$, the attention scores is calculated as:
$\bm{A_{\theta}}=\bm{\theta}_{l}\odot\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M})\right)$
where $\odot$ denotes element-wise multiplication. Since all elements of $\bm{\theta}$ are initially set to 1, this operation does not alter the attention scores.
The model then outputs a logit prediction $\bm{Y_{\theta}}$ for the next token, represented as a vector $\bm{Y_{\theta}}=[y_{1},y_{2},\dots,y_{\text{vocab}}]$. Our focus is on a specific logit $y_{\mathrm{gt}}$, associated with the correct label or ground truth token. For instance, the correct answer in a multiple-choice scenario or the initial token of the target sequence in a needle-in-a-haystack task.
We are particularly interested in the partial derivative $\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}$, which measures the sensitivity of the model’s output to changes in attention scores. We quantify the importance of a specific attention section by computing the mean of the derivative values within that section. Formally, we define the gradient-based importance as:
$\mathrm{Grad}(\bm{M},l)=\frac{\sum_{\bm{M}_{i,j}=1}{\frac{\partial y_{\mathrm{gt}}}{\partial\bm{\theta}_{l,i,j}}}}{\sum_{\bm{M}_{i,j}=1}{\bm{M}_{i,j}}}$
where $\bm{M}$ denotes the binary mask representing the targeted attention section, and $\mathrm{Grad}(\bm{M},l)$ indicates the average sensitivity of $y_{\mathrm{gt}}$ to changes within this section at layer $l$.
For comparison, we also define the average attention score of the section as:
$\mathrm{Att}(\bm{M},l)=\frac{\sum_{\bm{M}_{i,j}=1}{\bm{A}_{l,i,j}}}{\sum_{\bm{M}_{i,j}=1}{\bm{M}_{i,j}}}$
[…]
Empirically, we leverage a key-value retrieval task to analyze the importance of different sections of attention. The language model is tasked with retrieve the corresponding value, starting from its first generated token. $y_{gt}$ is set to be the first token of the correct value. The full task prompt is provided in Appendix A.1.
We calculate $\mathrm{Grad}(\bm{M},l)$ and $\mathrm{Att}(\bm{M},l)$ for both the middle Q-K section ($\bm{M}^{\mathrm{middle}}$) and the last Q-K section ($\bm{M}^{\mathrm{last}}$) across 100 randomly generated samples. The hyperparameters are set as follows: the number of sink tokens ($si$) is 64, the sliding window size ($sl$) is 128, and the length of the last attention section ($last$) is also set to 128. Each input sequence consists of approximately 2,000 tokens. Finally, we visualize the results for $\mathrm{Grad}(\bm{M},l)$ and $\mathrm{Att}(\bm{M},l)$ in Figure 4. For clarity, attention scores and gradient values are both summed across all heads within each layer.
2.4 TriangleMix
Methods Average
2Wiki
GovRep
Hotpot
LCC
MNews
MF-en
PsgCnt
PsgRtr
Qasper
Rbench
Samsum
TREC
TrvQA
Llama-3.1-8B-Instruct 54.8 48.1 34.4 61.6 66.6 25.6 56.9 16.8 99.7 44.9 52.7 42.6 71.0 91.6
Streaming 34.7 17.4 32.2 19.8 65.6 24.6 25.6 5.7 11.8 23.3 54.1 40.6 52.7 77.3
Triangle 47.0 40.6 33.2 52.3 63.6 24.6 53.5 7.9 42.0 46.3 46.5 43.0 67.0 90.1
StreamingMix 48.9 36.4 34.3 39.4 64.9 25.8 42.3 18.4 98.3 42.4 53.6 42.6 51.0 86.8
DuoAttention 54.5 44.6 34.1 59.1 68.0 25.3 55.0 15.7 99.7 43.9 58.8 42.7 71.3 91.0
[…]
Llama-3-8B-Instruct-262k 44.2 21.0 34.3 26.4 46.1 26.3 50.2 4.7 95.0 30.5 43.3 41.1 68.3 86.8
Streaming 30.8 16.8 34.3 18.5 41.8 25.3 39.6 0.3 10.0 23.4 36.5 38.6 49.0 66.0
Triangle 36.2 17.1 34.1 21.2 35.5 25.3 50.0 5.0 32.0 28.9 32.4 40.5 63.0 85.9
StreamingMix 39.9 18.5 34.4 19.8 42.3 25.9 47.4 5.3 85.0 30.0 38.3 40.1 48.3 83.3
DuoAttention 42.6 19.2 34.6 25.1 44.4 26.1 51.8 2.3 87.7 29.6 41.1 40.4 65.7 85.9
[…]
Qwen2.5-7B-Instruct 46.8 24.5 30.7 32.2 62.2 22.3 41.3 13.1 99.3 24.4 60.7 42.5 67.0 88.7
Streaming 28.3 10.1 31.3 9.5 61.7 21.5 20.7 6.5 8.2 11.4 44.8 40.2 51.7 49.9
Triangle 36.9 17.4 27.5 24.0 48.5 17.9 38.5 5.6 43.9 22.3 39.6 41.2 65.0 88.1
StreamingMix 45.2 24.4 31.9 38.5 68.2 22.5 43.5 10.7 93.0 24.8 60.9 42.0 43.7 83.1
[…]
Table 2: Comparison with static sparsity methods on Longbench. TriangleMix is our proposed method. StreamingMix is a hybrid baseline that uses dense attention in shallow layers but replaces Triangle attention with a streaming pattern in deeper layers. The number of dense layers matches that of TriangleMix.
From Figure 4, we observe that both the Middle Q-K and Last Q-K sections have low average attention scores, with the Middle Q-K section slightly higher across all layers. However, when examining the gradient values, we find that the derivatives in the Last Q-K section are significantly larger than those in the Middle Q-K section. This indicates that the model output is more sensitive to changes in the attention scores of the Last Q-K section compared to the Middle Q-K section.
Besides, for the last layer, we observe that we always have $\mathrm{Grad}(\bm{M},l)=0$. This phenomenon is expected due to the specific structure of the attention mechanism. In the last layer, the logit corresponding to the first predicted token is computed solely based on the last query vector and the entire key matrix $\bm{K}^{T}$. Formally, this corresponds to computing only the last row of the attention map, represented as $\mathrm{Softmax}(\bm{Q}_{\mathrm{last}}\bm{K}^{T})$. Since the gradient measure $\mathrm{Grad}(\bm{M},l)$ is defined based on the derivative of the ground truth output logit with respect to parameters associated with the attention sections defined by the mask, and because the first output token logit does not depend on the Middle Q-K sections in the last layer (which involve query vectors from other positions), the resulting gradient naturally becomes zero.
During the training phase, this observation does not apply. The training loss is typically computed using all output logits across the entire input sequence, which requires calculating the full attention matrix in the last layer. However, during inference, since our primary interest lies only in the prediction, we can optimize the computation in the final layer by restricting attention calculation to $\mathrm{Softmax}(\frac{1}{\sqrt{d}}\bm{Q}_{\mathrm{last}}\bm{K}^{T})$.
Moreover, we observe that the gradient values for the Middle Q-K sections follow a consistent trend across all models: they remain moderate in the shallow layers but drop sharply in deeper layers. As shown in Figure 2, this pattern is evident. For Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, the decline occurs between layers 10 and 20, while for Qwen2.5-7B-Instruct, it starts after layer 20.
This suggests that the attention from the Middle Q-K sections in deep layers contribute less to the output logits. Notably, the number of elements in the Middle Q-K sections scales quadratically with the input token length $N$, whereas the elements in the Streaming and Last Q-K sections grow only linearly with $N$. Therefore, if the computation of the Middle Q-K sections is skipped in deep layers, the time complexity of attention can be reduced from $O(N^{2})$ to $O(N)$. And we assume such skipping will not cause significant loss in output quality.
Based on these observations, we propose TriangleMix, an effective and efficient static attention pattern designed for long-context LLMs. We denote the starting layer of Triangle attention as $L_{\mathrm{tri\_start}}$. For layer $l\leq L_{\mathrm{tri\_start}}$, we apply standard dense attention:
$\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c(1-\bm{M})\right)$
where $\bm{M}$ is the standard causal attention mask.
In layer $l>L_{\mathrm{tri\_start}}$, we skip the Middle Q-K sections and apply a Triangle-shape attention pattern:
$\mathrm{Softmax}\left(\frac{1}{\sqrt{d}}\bm{QK}^{T}-c\left(1-(\bm{M}-\bm{M}^{\text{middle}})\right)\right)$
where $\bm{M}^{\text{middle}}$ represents the mask corresponding to the Middle Q-K sections.
Thus, in deep layers, the attention complexity becomes $O(N)$, and there is no need for dynamic estimation of the attention mask, making the computation highly efficient. Furthermore, we can optionally adopt dynamic sparse attention in shallow layers to further accelerate computation, while keeping the static Triangle pattern in deep layers.
We acknowledge that the Triangle-shape attention pattern was first introduced in Scbench Li et al. (2024a). However, Scbench applies the Triangle-shape attention uniformly across all layers, leading to notable performance degradation. In contrast, by using a gradient-based probing method, we find that the Middle Q-K sections remain important in shallow layers. Therefore, by combining dense (or dynamic sparse) attention in shallow layers and Triangle-shape attention in deep layers, TriangleMix preserves model performance while achieving significant efficiency gains.
3 Evaluations
3.1 Settings
LLMs. We evaluate our method on the following long-context language models. (1) Llama-3.1-8B-Instruct, a state-of-the-art model from Meta with a context window of up to 128K tokens Grattafiori et al. (2024). (2) Llama-3-8B-Instruct-262K, a fine-tuned variant from Llama-3-8B-Instruct that supports up to 262K tokens GradientAI (2024). (3) Qwen2.5-7B-Instruct, a model from Alibaba with a maximum context window of 128K Yang et al. (2024).
Static Sparsity Baselines. We compare our method with four static sparsity baselines. The first is the Streaming pattern Xiao et al. (2023), applied to all layers with $si=8$ and $sl=512$. The second is Triangle Attention applied to all layers, using $si=8$, $sl=512$, and $last=128$. The third is a hybrid baseline StremingMix, where dense attention is used for layers $l\leq L_{\text{tri\_start}}$, and streaming attention is used for deeper layers.
We also compare against DuoAttention, which learns a separate sparsity pattern for each attention head Xiao et al. (2024). DuoAttention provides a trained pattern for Llama-3.1-8B-Instruct; we train a sparsity pattern for Llama-3-8B-Instruct-262K using the authors’ script. However, for Qwen2.5-7B-Instruct, the training does not converge, so its result is omitted. At the prefilling stage, half of the heads follow the streaming pattern, while the other half use dense attention.
Dynamic Sparsity Baselines. We compare with two dynamic sparsity methods: MInference Jiang et al. (2024a) and FlexPrefill Lai et al. (2025). We apply default sparsity settings of all LLMs from MInference. For FlexPrefill, we test its hyperparameter $\gamma=0.90$ and $\gamma=0.95$. Additionally, our method can be combined with dynamic sparsity attention: Dynamic sparsity (MInference or FlexPrefill) is applied for layers $l\leq L_{\text{tri\_start}}$, while Triangle Attention is used for deeper layers. This combination reduces runtime overhead because Triangle Attention avoids runtime mask estimation in dynamic sparse attention. We use “Ours + MInference” and “Ours + FlexPrefill” to denote the corresponding combined versions.
Hyperparameters. The hyperparameter $L_{\text{tri\_start}}$ determines the layer from which the Triangle Attention pattern is applied. For layers $l\leq L_{\text{tri\_start}}$, we use dense attention or dynamic sparsity methods. For layers $l>L_{\text{tri\_start}}$, we switch to the static Triangle Attention. Based on the observations in Figure 4, we set $L_{\text{tri\_start}}=16$ for Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K, and $L_{\text{tri\_start}}=20$ for Qwen2.5-7B-Instruct. For the Triangle mask, we use: sink tokens $si=8$, sliding window size $sl=512$, and last window size $last=128$, which is the same with the sparse sparsity baselines.
Benchmarks. We evaluate on two benchmarks: (1) RULER: A synthetic long-context benchmark including tasks like hay-in-the-stack, KV retrieval, variable tracking, and QA. We evaluate at input lengths of 4K, 8K, 16K, 32K, 64K, and 128K, with 100 samples per task-length pair Hsieh et al. (2024). (2) LongBench: A realistic long-context benchmark with tasks such as single-doc and multi-doc QA, few-shot learning, code generation, retrieval, summarization, and counting Bai et al. (2023). We evaluate all English tasks.
Specifically, Qwen2.5-7B-Instruct adopts Yarn scaling to improve its long context ability for input length larger than 32K Peng et al. (2023). We report the performance of the origin Qwen2.5-7B-Instruct model for RULER length < 32K, and the Yarn-scaled model for RULER length 32K, 64K and 128K. For LongBench, we only evaluate the origin Qwen2.5-7B-Instruct model for simplicity.
3.2 Effectiveness of TriangleMix
Methods 4K 8K 16K 32K 64K 128K Avg.
Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
Streaming 64.1 55.4 40.5 28.9 26.7 3.3 36.5
Triangle 88.8 88.3 82.8 72.6 65.0 39.0 72.7
StreamingMix 94.3 91.8 90.2 86.2 79.2 71.9 85.6
DuoAttention 95.7 93.1 88.4 84.3 82.5 64.0 84.7
[…]
Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
Streaming 49.4 38.7 33.4 30.2 26.5 21.7 33.3
Triangle 88.1 84.7 80.6 71.0 65.0 55.4 74.1
StreamingMix 87.3 83.2 80.8 79.9 77.6 70.8 79.9
DuoAttention 92.8 91.9 89.0 85.0 81.1 76.1 86.0
[…]
Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
Streaming 55.6 47.8 35.9 30.6 28.1 21.0 36.5
Triangle 88.4 81.9 74.0 58.3 51.0 14.6 61.4
StreamingMix 93.3 90.4 89.4 81.7 76.4 63.6 82.5
[…]
Table 3: Comparison with static sparsity methods on RULER. Llama-3.1, Llama-3-262K, Qwen2.5 are abbreviations for Llama-3.1-8B-Instruct, Llama-3-8B-262K, and Qwen2.5-8B-Instruct. The same applies to Table 4.
Comparison with Static Sparsity. Table 3 and Table 2 present the evaluation results of TriangleMix and various static sparsity baselines. TriangleMix consistently outperforms all static baselines and remains nearly lossless performance across all input lengths. In contrast, other static methods usually suffer noticeable performance degradation. DuoAttention performs comparably to TriangleMix on LongBench using Llama-3.1-8B-Instruct. However, its performance is worse than TriangleMix on RULER, especially deteriorates at longer lengths (e.g., 64K and 128K). Furthermore, DuoAttention requires a separate training phase to learn head-wise sparsity patterns, which introduces additional computational overhead and may fail to converge on models such as Qwen-2.5-7B-Instruct.
On the LongBench benchmark, we observe that the full Triangle pattern performs poorly on PassageRetrieval tasks, while the StreamingMix pattern underperforms on the in-context learning task (TREC). These findings suggest that attention over the Middle Q-K section in shallow layers is crucial for retrieval tasks, while attention over the Last Q-K section in deeper layers plays a key role in supporting in-context learning.
Comparison with Dynamic Sparsity. Table 4 and Table 8 report results comparing TriangleMix with dynamic sparsity methods, as well as their combinations. FlexPrefill when $\gamma=0.90$ usually leads to performance degradation, and MInference generally outperforms it. While the performance of MInference and FlexPrefill with $\gamma=0.95$ is similar on the RULER benchmark, MInference still achieves better results on LongBench. When combined with either MInference or FlexPrefill, TriangleMix always fully retains the original performance of the dynamic sparsity methods. Moreover, combining TriangleMix with FlexPrefill even yields slight performance gains, particularly with Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct.
Overall, TriangleMix preserves the full accuracy of dense attention across tasks while significantly reducing time complexity from $O(N^{2})$ to $O(N)$ for layers $l>L_{\text{tri\_start}}$.
Methods 4K 8K 16K 32K 64K 128K Avg.
Llama-3.1 96.6 95.3 94.8 91.3 86.3 78.1 90.4
MInference 96.3 95.1 95.0 90.5 86.8 75.0 89.8
[…]
FlexPrefill (0.90) 90.1 89.7 89.4 89.1 82.0 74.1 85.7
[…]
FlexPrefill (0.95) 95.0 94.7 94.5 92.8 86.5 75.9 89.9
[…]
Llama-3-262k 93.4 90.3 88.8 85.1 82.2 79.4 86.5
MInference 93.4 90.7 88.8 85.1 82.6 80.1 86.8
[…]
FlexPrefill (0.90) 82.4 80.0 80.7 76.2 74.5 69.1 77.1
[…]
FlexPrefill (0.95) 90.3 87.7 88.0 83.8 80.2 75.7 84.3
[…]
Qwen2.5 95.8 93.6 92.6 84.5 81.9 67.4 86.0
MInference 95.6 93.8 92.8 84.9 79.1 68.0 85.7
[…]
FlexPrefill (0.90) 84.4 79.4 82.4 71.4 65.4 47.7 71.8
[…]
FlexPrefill (0.95) 91.1 89.7 88.5 75.4 72.5 51.2 78.1
[…]
Table 4: Comparison with dynamic sparsity methods on RULER.
3.3 Analysis of Start Layer
Figure 5: Average RULER score at 64K length for different $L_{\mathrm{tri\_start}}$ values.
In this section, we investigate the choice of $L_{\text{tri\_start}}$ for each LLM. We evaluate all models on the 64K length tasks in the RULER benchmark, sweeping $L_{\text{tri\_start}}$ from 0 to the total number of layers with a step size of 2. As shown in Figure 5, Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit similar patterns: setting $L_{\text{tri\_start}}=12$ retains 99.7% and 100.0% of their original performance, respectively. This indicates that 62.5% of the layers can adopt the $O(N)$ Triangle Attention without noticeable degradation. In contrast, Qwen-2.5-7B-Instruct requires a higher threshold: setting $L_{\text{tri\_start}}=20$ retains 97.9% of the performance, corresponding to a sparsity ratio of 28.6%.
Method 32K 64K 128K
Dense 44 179 750
MInference 94 (0.5x) 152 (1.2x) 229 (3.3x)
FlexPrefill (0.95) 32 (1.4x) 71  (2.5x) 220 (3.4x)
FlexPrefill (0.90) 24 (1.8x) 52  (3.4x) 143 (5.2x)
[…]
Table 5: Attention kernel latency (ms) for one layer measured on Llama-3.1-8B-Instruct.
Figure 6: Time-to-first-token (TTFT) in seconds measured on Llama-3.1-8B-Instruct.
Method 32K 48K 64K 80K 96K 112K 128K
Dense 4.1 7.3 11.2 15.9 21.3 27.5 34.5
MInference 5.5 (+34%) 7.8 (+7%) 10.1 (-10%) 12.3 (-23%) 13.4 (-37%) 15.9 (-42%) 18.0 (-48%)
FlexPrefill (0.95) 3.6 (-12%) 5.5 (-25%) 7.7 (-31%) 9.9 (-38%) 12.3 (-42%) 15.0 (-45%) 17.8 (-48%)
TriangleMix 3.6 (-12%) 5.9 (-19%) 8.6 (-23%) 11.7 (-26%) 15.2 (-29%) 19.1 (-31%) 23.4 (-32%)
Ours + MInference 4.2 (+2%) 6.0 (-18%) 7.7 (-31%) 9.5 (-40%) 10.9 (-49%) 12.7 (-54%) 14.5 (-58%)
Ours + FP (0.95) 3.4 (-17%) 5.2 (-29%) 7.2 (-36%) 9.2 (-42%) 11.3 (-47%) 13.6 (-51%) 16.1 (-53%)
FlexPrefill (0.90) 3.3 (-20%) 5.1 (-30%) 6.8 (-39%) 8.9 (-44%) 11.1 (-48%) 13.6 (-51%) 15.5 (-55%)
Ours + FP (0.90) 3.3 (-20%) 4.5 (-38%) 6.7 (-40%) 8.5 (-47%) 10.5 (-51%) 12.6 (-54%) 14.5 (-58%)
Table 6: Time-to-first-token (TTFT) in seconds measured on Llama-3.1-8B-Instruct. We separate FlexPrefill ($\gamma=0.90$) as it generally leads to accuracy degradation, whereas the other settings do not.
3.4 Efficiency of TriangleMix
We implement Triangle Attention using Triton Tillet et al. (2019), with further implementation details provided in Appendix A.2. All experiments are conducted on a single NVIDIA A100 80GB GPU. Dense attention is implemented using FlashAttention Kwon et al. (2023).
Attention Kernel Latency. We first benchmark the average attention kernel time per layer on Llama-3.1-8B-Instruct at sequence lengths of 32K, 64K, and 128K. As shown in Table 5, Triangle Attention significantly outperforms dense attention, achieving 3.7×3.7$\times$3.7 × to 15.3×15.3$\times$15.3 × speedup. The speedup mainly comes from Triangle Attention’s linear complexity ($O(N)$), which makes it much more efficient than $O(N^{2})$ dense attention. Besides, unlike dynamic sparsity methods, it doesn’t require estimating attention block indices during runtime, which saves additional overhead.
End-to-End TTFT. We further measure time-to-first-token (TTFT) across sequence lengths of 32K, 48K, 64K, 96K, 112K, and 128K. Results for Llama-3.1-8B-Instruct are reported in Table 6. TriangleMix applies dense attention in shallow layers and Triangle Attention in deeper layers, reducing TTFT by 12%–32%. It is also compatible with dynamic attention for enhanced performance. For example, integrating MInference with TriangleMix reduces TTFT from 18.0s to 14.5s (a 19% decrease). Under the lossless setting, TriangleMix combined with FlexPrefill ($\gamma=0.95$) achieves the fastest TTFT for 32K–80K sequences, while TriangleMix with MInference is optimal for 96K–128K. We visualize the results of some selected methods in Figure 6.
4 Discussion
In this section, we discuss why Middle Q-K attention can be safely removed. Our hypothesis is that the observed phenomena arise from a Train–Test Misalignment. During training, the loss is applied uniformly to all positions in the input sequence. For example, if the input contains 2048 tokens, the model is also trained to predict token 1001 given tokens 1–1000. However, during inference, the primary goal is to predict tokens after the prompt (such as token 2049), making predictions of tokens inside the prompt (such as token 1001) largely irrelevant.
To validate this assumption, we conduct an experiment on Llama-3.1-8B-Instruct. We sample 100 Wikipedia articles and extract the first 2048 tokens of each. In the last 16 layers, we apply Triangle Attention, skipping attention for Q indices between 512 and 1024. We then measure changes of the average perplexity (PPL) relative to dense attention in two output spans.
Method PPL (1024–1152) PPL (1920–2048)
Dense 8.02 7.79
TriangleMix 8.34 (+0.32) 7.81 (+0.02)
Table 7: With 2048 input tokens, we measure perplexity on two output spans. Triangle attention is applied, skipping attention between Q indices 512–1024 in the last 16 layers. Experiments are on Llama-3.1-8B-Instruct.
The results in Table 7 show that skipping attention of Q index 512–1024 in deep layers noticeably increases perplexity when predicting tokens in the 1024–1152 range. For these positions, the skipped attention effectively becomes the Last Q-K attention necessary for predicting the next token. In contrast, for positions beyond 1920, the effect is minimal.
The Middle Q-K region is important primarily for predicting tokens within the prompt, while for generation tasks focused on tokens after the prompt, the Middle Q-K in deeper layers is largely redundant. Although further theoretical work is needed, this observation suggests that selective sparsification of attention during inference can be done safely.
5 Related Works
Static Sparsity Attention. Early methods employ fixed sparse patterns, such as strided Child et al. (2019), dilated Ding et al. (2023), sliding window Jiang et al. (2023), and mixed patterns Beltagy et al. (2020), typically requiring training from scratch. DuoAttention Xiao et al. (2024) introduces head-level static sparsity but needs additional offline training. Training-free approaches like Streaming Xiao et al. (2023) and Triangle Li et al. (2024a) attention often degrade accuracy for long contexts. In contrast, the proposed TriangleMix attention pattern significantly mitigates performance loss and is nearly lossless in accuracy.
Dynamic Sparsity Attention. Methods like MInference Jiang et al. (2024a), FlexPrefill Lai et al. (2025), and XAttention Xu et al. (2025) dynamically determine attention masks online, maintaining high accuracy but incurring computational overhead for moderately long contexts. TriangleMix can enhance these dynamic approaches for further acceleration .
Long-context LLM Inference. FlashAttention Dao et al. (2022) speeds up attention by reducing memory access through fused operations. PagedAttention Kwon et al. (2023) improves decoding by managing KV cache allocation efficiently. KV cache optimizations techniques also include token-level eviction (SnapKV Li et al. (2024b)) and query-aware cache selection (Quest Tang et al. (2024)). TriangleMix is orthogonal to these approaches.
6 Conclusion
In this paper, we propose a hybrid attention pattern for long-context LLMs, where dense attention is used in shallow layers and Triangle attention is applied in deeper layers. This pattern is training-free and achieves nearly lossless accuracy. Our method achieves a 12% to 32% time-to-first-token reduction across 32K to 128K inputs. Moreover, it can be further accelerated when combined with dynamic sparsity attention.
Limitations
One limitation of our method is that the sparsity ratio is fixed within the pretrained model. We observe that Llama-3.1-8B-Instruct and Llama-3-8B-Instruct-262K exhibit higher inherent sparsity compared to Qwen2.5-7B-Instruct, which constrains the achievable acceleration on the latter. It may be possible to enhance the sparsity of models like Qwen2.5-7B-Instruct through a post-training optimization process, which we leave for future exploration.
Appendix A Appendix
A.1 Task to Probe Attention Importance
We use a key-value retrieval task to analyze the importance of different attention sections. The language model is prompted to retrieve the correct value, beginning from its first generated token. The prompt is as follows:
[…]
We randomly generate pairs of keys and values using Version 4 UUID strings.
A.2 Implementation of Triangle Attention
We implement Triangle Attention using Triton Tillet et al. (2019), with details provided in Algorithm 1. For each row index, we apply a different attention mechanism: if the row index is less than $N-N_{\text{last}}$, we apply Streaming Attention; otherwise, we apply chunked FlashAttention to increase GPU utilization. The outputs from both segments are then merged to produce the final result.
Algorithm 1 Fused Triangle Attention
Input data: $\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V}\in\mathbb{R}^{N\times d_{h}}$
Input triangle shape: sink token number $N_{\mathrm{sink}}$, sliding window size $N_{\mathrm{window}}$, last rows number $N_{\mathrm{last}}$
Input kernel block shape: $B_{M},B_{N}$
 Calculate best number of splits $S$ for last row
 Initialize $\boldsymbol{O}\leftarrow(0)^{N\times d_{h}}$
 Initialize output buffer $\boldsymbol{O}_{\mathrm{last}}\leftarrow(0)^{SN_{\mathrm{last}}\times d_{h}}$
 Initialize LSE buffer $\boldsymbol{l}_{\mathrm{last}}\leftarrow(-\inf)^{SN_{\mathrm{last}}}$
Parallelized in GPU
for $i\leftarrow 1$ to $\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil+S\lceil\frac{N_{\mathrm{last}}}{B_{M}}\rceil$ do
  Initialize $\boldsymbol{O}_{\text{chip}}\leftarrow(0)^{B_{M}\times d_{h}}$
  Initialize $\boldsymbol{m}\leftarrow(-\inf)^{B_{N}}$
  Initialize $\boldsymbol{s}\leftarrow(0)^{B_{M}}$
  if $i\leq\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil$ then
    Upper part: the same as streaming attention
   Load $\boldsymbol{Q}_{\text{chip}}\leftarrow\boldsymbol{Q}^{iB_{M}:(i+1)B_{M}}$
    Loop through sink tokens
   for $j\leftarrow 1$ to $\lceil\frac{N_{\mathrm{sink}}}{B_{N}}\rceil$ do
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$
    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
    $\mathrm{flash\_attn}(\boldsymbol{Q}_{\text{chip}},\boldsymbol{K}_{\text{chip}},\boldsymbol{V}_{\text{chip}},\boldsymbol{O}_{\text{chip}},\boldsymbol{m},\boldsymbol{s})$
   end for
    Loop through sliding window
   for $j\leftarrow\lceil\frac{N-N_{\mathrm{window}}}{B_{N}}\rceil$ to $\lceil\frac{N}{B_{N}}\rceil$ do
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}\in\mathbb{R}^{B\times d_{h}}$
    $\mathrm{flash\_attn}(\boldsymbol{Q}_{\text{chip}},\boldsymbol{K}_{\text{chip}},\boldsymbol{V}_{\text{chip}},\boldsymbol{O}_{\text{chip}},\boldsymbol{m},\boldsymbol{s})$
   end for
    Write outputs
   Save $\boldsymbol{O}^{iB_{M}:(i+1)B_{M}}\leftarrow\boldsymbol{O}_{\text{chip}}$
  else
    Last rows: split in to $S$ chunks
   Chunk index $c\leftarrow\lfloor(i-\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil)/{\lceil\frac{N_{\mathrm{last}}}{B_{M}}\rceil}\rfloor$
   Chunk offset $b\leftarrow(i-\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil)\mod{\lceil\frac{N_{\mathrm{last}}}{B_{M}}\rceil}$
   Q index $i_{Q}\leftarrow\lceil\frac{N-N_{\mathrm{last}}}{B_{M}}\rceil+b$
   Load $\boldsymbol{Q}_{\text{chip}}\leftarrow\boldsymbol{Q}^{i_{Q}B_{M}:(i_{Q}+1)B_{M}}$
   for $j\leftarrow c\lceil\frac{N}{SB_{N}}\rceil$ to $(c+1)\lceil\frac{N}{SB_{N}}\rceil$ do
    Load $\boldsymbol{K}_{\text{chip}}\leftarrow\boldsymbol{K}^{jB_{N}:(j+1)B_{N}}$
    Load $\boldsymbol{V}_{\text{chip}}\leftarrow\boldsymbol{V}^{jB_{N}:(j+1)B_{N}}$
    $\mathrm{flash\_attn}(\boldsymbol{Q}_{\text{chip}},\boldsymbol{K}_{\text{chip}},\boldsymbol{V}_{\text{chip}},\boldsymbol{O}_{\text{chip}},\boldsymbol{m},\boldsymbol{s})$
   end for
    Write outputs
   Save $\boldsymbol{O}_{\mathrm{last}}^{(i-i_{Q})B_{M}:(i-i_{Q}+1)B_{M}}\leftarrow\boldsymbol{O}_{\text{chip}}$
   Save $\boldsymbol{l}_{\mathrm{last}}^{(i-i_{Q})B_{M}:(i-i_{Q}+1)B_{M}}\leftarrow\ln\boldsymbol{s}+\boldsymbol{m}$
  end
end for
Merge last row output buffer
$\boldsymbol{O}^{(N-N_{\mathrm{last}}):N}\leftarrow\mathrm{merge\_output}(\boldsymbol{O}_{\mathrm{last}},\boldsymbol{l}_{\mathrm{last}})$
A.3 Additional Longbench Data
Table 8 compares our method with dynamic sparsity attention on Longbench. When combined with MInference, our method achieves similar performance to MInference alone. Integrating with FlexPrefill yields a slight improvement over its original results.
Methods Average
2Wiki
GovRep
Hotpot
LCC
MNews
MF-en
PsgCnt
PsgRtr
Qasper
Rbench
Samsum
TREC
TrvQA
Llama-3.1-8B-Instruct 54.8 48.1 34.4 61.6 66.6 25.6 56.9 16.8 99.7 44.9 52.7 42.6 71.0 91.6
MInference 54.9 48.1 34.5 61.4 67.0 25.9 57.0 17.5 99.7 44.7 52.5 42.6 71.3 91.9
[…]
FlexPrefill; $\gamma=0.90$ 44.0 31.0 32.8 48.8 59.1 25.1 51.3 4.0 39.0 37.1 50.8 41.5 65.3 86.2
[…]
FlexPrefill; $\gamma=0.95$ 48.4 39.4 33.7 58.3 67.2 25.6 55.6 4.0 47.7 42.5 53.8 41.9 69.7 90.2
[…]
Llama-3-8B-Instruct-262k 44.2 21.0 34.3 26.4 46.1 26.3 50.2 4.7 95.0 30.5 43.3 41.1 68.3 86.8
MInference 44.1 20.5 34.3 25.8 46.2 26.3 50.2 4.3 95.0 30.7 44.7 40.5 68.3 86.8
[…]
FlexPrefill; $\gamma=0.90$ 33.6 16.8 32.7 23.0 46.8 25.6 45.6 2.3 8.0 23.2 41.4 36.8 61.0 73.3
[…]
FlexPrefill; $\gamma=0.95$ 35.0 17.7 33.0 25.1 36.8 26.1 49.1 3.3 13.7 28.3 36.2 39.1 65.0 81.7
[…]
Qwen2.5-7B-Instruct 46.8 24.5 30.7 32.2 62.2 22.3 41.3 13.1 99.3 24.4 60.7 42.5 67.0 88.7
MInference 46.8 24.5 30.7 32.7 61.6 22.3 40.8 12.8 100.0 24.3 60.2 42.1 68.0 89.3
[…]
FlexPrefill; $\gamma=0.90$ 37.3 17.5 29.3 27.6 64.0 21.8 39.0 7.0 28.7 15.1 51.4 39.3 64.7 79.3
[…]
FlexPrefill; $\gamma=0.95$ 40.8 20.2 29.7 33.9 66.8 21.8 40.4 4.0 48.7 17.1 54.8 40.5 67.3 85.5
[…]
Table 8: Comparison with dynamic sparsity methods on LongBench.